For the first 3 questions, refer to the following

Problems 1-3, which ask you to explore expressions for the interatomic bonding energy, are confusing as you need the answer to problem 3 to answer problem 2 - they are a bit inverted. Please work you way through problems 1 and 2, and the answer to problem 3 is already contained in the earlier problems. Sorry for the confusion!
Answer #1

Given EN= -A/r + B/rn

Hence d(EN/dr) =A/r2 – nB/r(n+1) = 0 for minimum energy, r becomes r0 ( the minium interatomic spacing

A/ro2 = nB/ro(n+1)

ro (n+1)/ro2= nB/A

ro(n+1-2)= (nB/A )

ro= (nB/A )1/(n-1)

EN( Minimum)= -A/(nB/A)1/(n-1) +B/(nB/A)n/(n-1)